OpenStudy (anonymous):

Use cross products to the solve proportion. 1. c/22 = 5/11 2. 1/6 = 8/t

OpenStudy (idku):

Do you mean cross multiplication?

OpenStudy (anonymous):

Yeah.

OpenStudy (idku):

OK, so I'll give you an example and then you will do your problems, OK?

OpenStudy (anonymous):

Um, okay.

OpenStudy (idku):

\[\frac{3}{4}=\frac{x}{16}\] you cross multiply so lets take this literally "cross multiply", multiply diagonally. \[3 \times 16= 4 \times x\] \[48 = 4x\] divide both sides by 4 to eliminate/solve for x. \[12=x\] can you do you problems using this example?

OpenStudy (anonymous):

Um okay, I guess.. Thanks for helping.

OpenStudy (idku):

Anytime, just tell me what you get for both of them, to make sure you are doing it right.